How can I translate this word problem into an equation? Jim is 5 years older than Joe. If the sum of their ages is 55, how old are they?
Answer by nerdybill(7384)

How can I translate this word problem into an equation? Jim is 5 years older than Joe. If the sum of their ages is 55, how old are they?
.
Let x = age of Joe
then
x+5 = age of Jim
.
x + x+5 = 55
2x + 5 = 55
2x = 50
x = 25 years old (Joe)
.
Jim:
x+5 = 25+5 = 30 years old